Please help me find an equation of the line through (5,-3) and (10,7)
Answer by rfer(16322)

m=(7--3)/(10-5)
m=10/5
m=2
------------
y=mx+b
7=2*10+b
7=10+b
b=-3
-------------
y=2x-3